OpenStudy (anonymous):

Match each with its equivalent. a^1 A.a^n a^0 B. =a a^-n C.=1/a^2 1/a^-2 D.=1

OpenStudy (anonymous):

a^1 = B a^0 = D a^-n = 1/a^n 1 /a^-2 = a^2

OpenStudy (nilankshi):

good~